# references.bib

### v1

@@ 5 Conclusion
This study analyzes CBA, an emerging approach with considerable potential to address the competency gaps inherent in traditional self-service analytics. Through the use of LLMs, CBA facilitates natural language interactions, enabling end users to independently perform tasks such as data retrieval, preparation, and insight generation. In this context, CBA represents a new form of task delegation, shifting the responsibility for generating and interpreting information from human experts to AI, thus making advanced analytics more accessible to users without technical expertise. A focus of this study is Text-to-SQL, a well-established semantic parsing technology that has seen transformative advancements through the application of LLMs, enhancing its ability to translate natural language requests into structured SQL statements.
@@
The models developed in this study examine the conditions under delegation to AI, through PS and FS, outperforms the delegation of tasks to human experts. PS is effective when the accuracy ($\alpha$) exceeds a specified threshold, as outlined in condition (1). FS, however, is only beneficial when three conditions are met: the AI delegation condition ((2)), FS feasibility condition ((3)), and validation dominance condition ((4)).
@@
The study highlights the critical role of validation effectiveness ($\beta$) in FS. While FS can significantly improve decision-making when validation is robust, the use of user-based validation introduces risks, particularly when users lack the necessary expertise or when explanations are unclear. These risks can undermine decision quality, making FS less effective than PS and, in worst case, even less effective than delegation to a data engineer. This underscores the necessity of robust validation mechanisms to assist users in accurately evaluating AI-generated information. Without such mechanisms, the advantages of CBA may be significantly compromised.
This study also opens several avenues for future research. The uncertainties associated with user-based validation warrant further investigation. Future research could explore techniques to support end users in the validation process, including training programs to build technical expertise and AI-driven tools that provide guidance during validation. Moreover, there is an urgent need for empirical studies that more clearly define the boundaries of user-based validation based on explanations, particularly in real-world scenarios involving complex SQL queries and data models.
While the primary goal of CBA is to overcome the technical barriers that hinder the use of traditional self-service analytics, user-based validation poses a paradox. The process relies on the same competencies that were insufficient for conducting independent analyses in the first place. This raises a critical question: is it advisable to involve end users with limited technical expertise in the validation of results, given that they may face similar challenges in assessing the quality of the analysis as they did in performing it?
This question highlights the importance of more clearly delineating the boundaries of user-based validation. While CBA relies on end-user interaction, additional support mechanisms may be necessary to help users validate the accuracy of results without risking misinterpretation or circular reasoning. One promising avenue for future research is the development of hybrid approaches that combine expertise with automated validation mechanisms.
@@
Future studies could explore the implementation of automated techniques to independently assess the reliability of generated KPIs. Potential approaches might include statistical methods such as confidence intervals to evaluate KPI correctness, automated tools for schema alignment and comparison, and cross-validation with pre-validated KPIs. Integrating such methods could reduce reliance on user expertise, thereby enhancing the robustness and scalability of CBA systems.
@@
The models presented in this study rely on several simplifying assumptions. First, the gain from correct decisions is fixed at $+1$ and the loss from incorrect decisions at $-1$, which may not fully capture the variability observed in real-world scenarios. Additionally, the analysis adopts a static perspective, assuming a single execution cycle for KPI generation and validation, thus overlooking the iterative nature of real-world analytics. The models also omit costs associated with LLM usage, data storage, and processing, which may lead to an overestimation of the economic feasibility of CBA. Expanding these models to incorporate dynamic, cost-sensitive, and multi-stage decision processes could provide deeper insights.
In conclusion, this study illustrates the transformative potential of CBA in addressing the limitations of traditional self-service analytics. However, realizing its full potential depends on addressing the inherent challenges of validation, advancing user support mechanisms, and exploring automated alternatives. Through continued innovation, CBA has the potential to evolve into a reliable and indispensable tool for data-driven decision-making across a wide range of business contexts.
[…]

### v2

@@ 5 Conclusion
This study analyzes CBA, an emerging approach with considerable potential to address the competency gaps inherent in traditional self-service analytics. Through the use of LLMs, CBA facilitates natural language interactions, enabling end users to independently perform tasks such as data retrieval, preparation, and insight generation. In this context, CBA represents a novel approach to task delegation in business analytics, shifting the process of information generation from human experts to AI, thereby making advanced analytics more timely and accessible to a broader range of organizational roles. A focus of this study is Text-to-SQL, a well-established semantic parsing technology that has seen transformative advancements through the application of LLMs, enhancing its ability to translate natural language requests into structured SQL statements.
@@
The theoretical models developed in this study examine the conditions under delegation to AI, through PS and FS, outperforms the delegation of tasks to human experts. PS is effective when the accuracy ($\alpha$) exceeds a specified threshold, as outlined in condition (1). FS, however, is only beneficial when three conditions are met: the AI delegation condition (2), FS feasibility condition (3), and validation dominance condition (4).
@@
The study highlights the critical importance of validation effectiveness ($\beta$) in the context of FS. FS can exert a "boosting effect" on information generation, particularly when validation is robust. Even with low accuracy ( $\alpha$), a positive benefit can be derived from the delegation process by preventing incorrect information from being incorporated into operational decision-making. However, the reliance on user-based validation introduces significant uncertainties, especially when users lack the necessary expertise or when SQL queries and data models are characterized by high complexity. These challenges can have a ”devastating effect” on the quality of information generation, potentially making FS less effective than PS and, in the worst-case scenario, even less effective than delegation to a data engineer. This underscores the need for robust validation mechanisms that support in correctly evaluating AI-generated information. Without such mechanisms, the advantages of CBA – particularly when the accuracy ($\alpha$ ) of translating natural language requests into correct information is high – could be significantly diminished.
@@
A promising avenue for future research is the development of hybrid approaches that combine existing techniques (e.g., decomposition, visualization and dialogue techniques) to support code comprehension with automated validation techniques. Future studies could explore the implementation of automated methods for user-independent assessment of the reliability of generated KPIs. Potential approaches might include statistical methods, such as confidence intervals to evaluate the correctness of KPIs, and cross-validation with previously validated reports and KPIs. Integrating such methods could mitigate the risk of misclassification due to a lack of expertise, thereby enhancing the robustness of CBA.
@@
The theoretical models presented in this study are based on several simplifying assumptions. First, the gain from making the correct decision is fixed at $+1$, and the loss from making the wrong decision is set to $-1$. While these values determine the derived conditions, they may not fully capture the variability observed in real-world scenarios. Furthermore, the analysis adopts a static perspective, assuming a single execution cycle for KPI creation and validation, thereby overlooking the iterative nature of CBA. The models also disregard the costs associated with using LLMs, data storage, and processing, potentially leading to an overestimation of the economic feasibility of CBA. Additionally, they assume perfect data quality, which may not align with practical conditions. Extending these models to incorporate dynamic, cost-sensitive, multi-stage decision-making processes and data quality considerations could yield deeper and more realistic insights.
In summary, this study highlights the transformative potential of CBA in overcoming the limitations of traditional self-service analytics and advancing the capabilities of business analytics. However, to fully realize this potential, it is crucial to address inherent challenges related to accuracy and validation, while exploring the added value of automated validation techniques. Integrating CBA into business analytics can improve data-driven decision-making and enhance accessibility to advanced analytics. With continued innovation, CBA holds the promise of becoming a trusted and indispensable approach for informed decision-making across diverse business contexts.